OpenStudy (anonymous):

solve. 3sq(72) - 2sq(8) thank you!

OpenStudy (anonymous):

\[3\sqrt{2*36}-2\sqrt{4*2}=3*6\sqrt{2}-2*2\sqrt{2}=18\sqrt{2}-4\sqrt{2}=14\sqrt{2}\]

OpenStudy (anonymous):

\[14 \sqrt{2} \]

OpenStudy (anonymous):

decimal = 19.7989898732233

OpenStudy (anonymous):

thank you!